# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ 6.1 Proof of Theorem 8
The new technical ingredients in our proof are the large deviation inequalities for sparse $\alpha$-subexponential random variables. We first note that the large deviation inequalities in Theorem 2 and Corollary 1 can be extended to complex coefficients or complex matrices by treating the real and imaginary parts separately. For the convenience of our subsequent analysis, we summarize the necessary results below.
Let $a=(a_{1},\cdots,a_{n})^{T}\in\mathbb{C}^{n}$ and $A=(a_{ij})_{1\leq i,j\leq n}\in\mathbb{C}^{n\times n}$, where $A$ is diagonal-free and $A=A^{*}$. Consider a random vector $X=(X_{1},\cdots,X_{n})^{T}\in\mathbb{R}^{n}$ with independent centered $\alpha$-subexponential random components. Each $X_{i}=x_{i}\xi_{i}$, where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ has mean 0, variance 1, and satisfies $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for all $r\geq 1$. Then by Theorem 2, Theorem 3, and Corollary 1, there is a constant $C_{\alpha}>0$ such that for any $r\geq 2$,
@@
Without loss of generality, we assume $C_{\alpha}\geq 1$.
@@
For the third term on the right-hand side of (6.15), we use the nesting property of conditional $L_{r}$ norms to derive
@@
The proof of Proposition 1 follows the standard stochastic continuity argument as seen in [10, 24]. For completeness, we include the proof in Appendix D.
The proof of Theorem 12 follows from Proposition 1, the Lipschitz continuity of $G_{ij}$ and $m$, $\zeta$ is a decreasing function of $\eta$, and $\frac{1}{(n\eta)^{2}}\leq\zeta t$, and a union bound on the lattice set, following the same arguments as in [24]. We omit the details.
@@ 6.2 Proof of Theorem 11
Proof of Theorem 11.
Note that $\|BX\|^{2}=X^{T}B^{T}BX$. To apply Corollary 2 to matrix $A=B^{T}B$ and random vector $X$, note that $(B^{T}B)_{ij}=\langle B_{i},B_{j}\rangle$ and the expectation can be computed as
@@
Hence, applying the tail bound in Corollary 2 with
@@
$+s^{\max\{\frac{2}{\alpha},2\}}\cdot K^{2}\|B^{T}\|_{2,\infty}^{2},$
@@ Acknowledgments
Appendix A Auxiliary inequalities
@@
Appendix B Generalization to noncentered random variables
In this section, we briefly discuss the extension of our results to the concentration of quadratic forms of noncentered random variables. Let $Y_{1},\cdots,Y_{n}$ be independent random variables, and consider a symmetric, diagonal-free matrix $A$ for simplicity.
Denote $X_{i}=Y_{i}-\mathbb{E}(Y_{i})$. We rewrite
$Y^{T}AY=\sum_{i\neq j}a_{ij}Y_{i}Y_{j}=\sum_{i\neq j}a_{ij}(X_{i}+\mathbb{E}(Y_{i}))(X_{j}+\mathbb{E}(Y_{j})).$ (B.1)
Through straightforward calculation, we obtain
The concentration inequality for $Y^{T}AY-\mathbb{E}(Y^{T}AY)$ can be obtained by considering those of $X^{T}AX$ and $\sum_{i}b_{i}X_{i}$ separately. Assume the moments of $Y_{i}$ satisfy $\mathbb{E}|Y_{i}|^{r}\leq p_{i}(Kr)^{r}$ for $r\geq 1$. Then from the inequality $\|X_{i}\|_{L_{r}}=\|Y_{i}-\mathbb{E}(Y_{i})\|_{L_{r}}\leq 2\|Y_{i}\|_{L_{r}}$, we deduce that the $X_{i}$’s are independent centered random variables satisfying $\mathbb{E}|X_{i}|^{r}\leq p_{i}(2Kr)^{r}$ for $r\geq 2$. Consequently, the conclusion of Theorem 1 holds for $X^{T}AX$. Moreover, $\sum_{i=1}^{n}b_{i}X_{i}$ is a sum of independent random variables. Applying Corollary 1 with $\alpha=2$, we find that the tail bound for $\sum_{i=1}^{n}b_{i}X_{i}$ is a mixture of sub-gaussian and sub-exponential tails. For the sub-exponential tail, we can use the estimate that
Based on the preceding discussion, we now present a generalization of Theorem 1 for $\sum_{i\neq j}a_{ij}Y_{i}Y_{j}$, where the random variables $Y_{i}$ are not necessarily centered.
THEOREM 13.
Let $Y=(Y_{1},\cdots,Y_{n})^{T}$ be a random vector with independent sparse sub-exponential entries, i.e., each $Y_{i}$ satisfies $\mathbb{E}|Y_{i}|^{r}\leq p_{i}(Kr)^{r}$ for all $r\geq 1$, where $p_{i}\in(0,1)$ and $K>0$. Let $A$ be an $n\times n$ diagonal-free matrix. Then there exists an absolute constant $C>0$ such that for any $t>0$,
$\mathbb{P}\left(\left|Y^{T}AY-\mathbb{E}(Y^{T}AY)\right|\geq t\right)$ (B.2)
$\leq C^{\prime}\exp\left(-C\min\left\{\frac{t^{2}}{K^{4}(\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}+\sum_{i}b_{i}^{2}p_{i})},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{1/2}\right\}\right),$ (B.3)
where $b_{i}:=\max\{\sum_{j\neq i}a_{ij}\mathbb{E}(Y_{j}),\sum_{j\neq i}a_{ji}\mathbb{E}(Y_{j})\}$ and $\gamma_{1,\infty}$ is defined in (1.3).
Appendix C Proof of Theorem 10
@@
LEMMA C.1.
There is an absolute constant $C>0$ depending such that for any $D>0$, if $p\geq\frac{K^{2}e^{2(D+1)}(\log n)^{1+\frac{2}{\alpha}}}{n}$, with probability at least $1-2n^{-D}$,
$\|\widetilde{H}\|\leq C(1+\sqrt{D})\sqrt{np}.$ (C.1)
@@
Since the entries of $\widetilde{H}$ satisfy $\mathbb{E}|\widetilde{H}_{ij}|^{r}\leq K^{r}r^{\frac{r}{\alpha}}$ for all $r\geq 1$, applying Markov’s inequality with a union bound yields
$\mathbb{P}\left(|\widetilde{H}_{ij}|>2e^{D+1}K(\log n)^{\frac{1}{\alpha}}\right)\leq n^{-2-D},$ (C.2)
$\mathbb{P}\left(\|\widetilde{H}\|_{\infty}>2e^{D+1}K(\log n)^{\frac{1}{\alpha}}\right)\leq n^{-D}.$ (C.3)
Denote $B:=2e^{D+1}K(\log n)^{\frac{1}{\alpha}}$. Define $\check{H}_{ij}=\widetilde{H}_{ij}\mathbf{1}\{|\widetilde{H}_{ij}|\leq B\}$. Then $|\check{H}_{ij}-\mathbb{E}\check{H}_{ij}|\leq 2B$. [30, Theorem 4.8] implies
$\|\check{H}-\mathbb{E}\check{H}\|\leq C\left(\sqrt{np}+B\sqrt{\log n}+t\right)$ (C.4)
@@
$\|\mathbb{E}\check{H}\|$ $=n|\mathbb{E}\check{H}_{ij}|=n|\mathbb{E}\check{H}_{ij}-\mathbb{E}\widetilde{H}_{ij}|=n|\mathbb{E}\widetilde{H}_{ij}\mathbf{1}\{|\widetilde{H}_{ij}|>B\}|$ (C.5)
$\leq n\sqrt{\mathbb{E}\widetilde{H}_{ij}^{2}}\sqrt{\mathbb{P}(|\widetilde{H}_{ij}|>B)}\leq Cn\sqrt{p}n^{-1-D/2}\leq C\sqrt{np}.$ (C.6)
@@
$\|\widetilde{H}\|=\|\check{H}\|$ $\leq\|\check{H}-\mathbb{E}\check{H}\|+\|\mathbb{E}\check{H}\|$ (C.7)
$\leq C(1+\sqrt{D})(\sqrt{np}+Ke^{D+1}(\log n)^{\frac{1}{2}+\frac{1}{\alpha}})$ (C.8)
$\leq 2C(1+\sqrt{D})\sqrt{np},$ (C.9)
where in the last inequality, we use the assumption that $p\geq\frac{K^{2}e^{2(D+1)}(\log n)^{1+2/\alpha}}{n}$. ∎
We work on the conclusion of Lemma C.1. This result implies that for each eigenvalue $\lambda_{i}$ of $H$, $z_{i}:=\lambda_{i}+\mathrm{i}\eta$, where $\eta=LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, lies in a spectral domain
@@
$\mathbb{P}\left(\max_{i,j}\left|G_{ij}(z)-m(z)\delta_{ij}\right|\leq\delta\right)\geq 1-n^{-D}.$ (C.10)
@@
Therefore, taking $\delta=1/2$ in (C.10) and using the bound $|m|\leq 1$, we have
$\frac{3}{2}\geq\operatorname{Im}G_{kk}(z_{i})=\sum_{j=1}^{n}\frac{\eta}{(\lambda_{j}-\lambda_{i})^{2}+\eta^{2}}|u_{j}(k)|^{2}\geq\frac{1}{\eta}|u_{i}(k)|^{2},$ (C.11)
@@
Appendix D Proof of Proposition 1

### v3

@@ 7.1 Proof of Theorem 6
The new technical ingredients in our proof are the large deviation inequalities for sparse $\alpha$-subexponential random variables. We first note that the large deviation inequalities in Theorem 1 and Theorem 2 can be extended to complex coefficients or complex matrices by treating the real and imaginary parts separately. For the convenience of our subsequent analysis, we summarize the necessary results below.
Let $a=(a_{1},\cdots,a_{n})^{T}\in\mathbb{C}^{n}$ and $A=(a_{ij})_{1\leq i,j\leq n}\in\mathbb{C}^{n\times n}$, where $A$ is diagonal-free and $A=A^{*}$. Consider a random vector $X=(X_{1},\cdots,X_{n})^{T}\in\mathbb{R}^{n}$ with independent centered $\alpha$-subexponential random components. Each $X_{i}=x_{i}\xi_{i}$, where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ has mean 0, variance 1, and satisfies $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for all $r\geq 1$. Then by Theorem 1 and Theorem 2, there is a constant $C_{\alpha}>0$ such that for any $r\geq 2$,
@@
To derive (7.2), we first observe that $\|X_{i}^{2}-\mathbb{E}X_{i}^{2}\|_{L_{r}}\leq 2\|X_{i}^{2}\|_{L_{r}}\leq(p_{i}2^{r}K^{2r}r^{\frac{2}{\alpha}})^{1/r}$. Then we apply Theorem 2 to the centered random variables $X_{i}^{2}-\mathbb{E}X_{i}^{2}$ to obtain (7.2). Without loss of generality, we assume $C_{\alpha}\geq 1$.
@@
For the fourth term on the right-hand side of (7.16), we use the nesting property of conditional $L_{r}$ norms to derive
@@
The proof of Proposition 7.1 follows the standard stochastic continuity argument as seen in [13, 34]. For completeness, we include the proof in Appendix C.
The proof of Theorem 11 follows from Proposition 7.1, the Lipschitz continuity of $G_{ij}$ and $m$, $\zeta$ is a decreasing function of $\eta$, and $\frac{1}{(n\eta)^{2}}\leq\zeta t$, and a union bound on the lattice set, following the same arguments as in [34]. We omit the details.
@@ 7.2 Proof of Theorem 10
Note that $\|BX\|^{2}=X^{T}B^{T}BX$. To apply Theorem 1 to matrix $A=B^{T}B$ and random vector $X$, note that $(B^{T}B)_{ij}=\langle B_{i},B_{j}\rangle$ and the expectation can be computed as
@@
Hence, applying the tail bound in Theorem 1 with
@@
$+s^{\frac{2}{\alpha}}\cdot K^{2}\|B^{T}\|_{2,\infty}^{2},$
@@ Acknowledgments
Appendix A Auxiliary results
A.1 Auxiliary lemmas
@@
A.2 The Lambert W function
For real numbers $x$ and $y$, the equation
$ye^{y}=x$
has real solutions for $y$ only when $x\geq-e^{-1}$. When $x\geq 0$, there exists a unique solution given by the principal branch of the Lambert W function, denoted as $y=W_{0}(x)$. For $-e^{-1}\leq x<0$, there are exactly two solutions: the principal branch $y=W_{0}(x)$ and the lower branch $y=W_{-1}(x)$. The two branches of the Lambert W function are illustrated in Figure 1. We include a result about the Lambert W function that is used in our proof.
Figure 1: The blue curve is the principal branch $W_{0}(x)$ and the red curve is the lower branch $W_{-1}(x)$.
PROPOSITION A. 1.
For positive $u$ and $L$, the inequality $e^{u}\leq Lu$ is solvable only when $L\geq e$ and the solution is
$-W_{0}(-L^{-1})\leq u\leq-W_{-1}(-L^{-1}).$
Proof.
The inequality $e^{u}\leq Lu$ is equivalent to
$(-u)e^{-u}\leq-L^{-1}.$ (A.2)
This inequality (A.2) is solvable only when $-L^{-1}\geq-e^{-1}$ since the minimum of $xe^{x}$ is $-e^{-1}$. Specially, if $L=e$, then $u=1$. For $L\geq e$, $-L^{-1}\in[-e^{-1},0)$ and from the definition and behavior of Lambert W function (see Figure 1), we find that $W_{-1}(-L^{-1})\leq-u\leq W_{0}(-L^{-1})$. This completes the proof. ∎
Appendix B Proofs of Theorem 8 and Theorem 9
B.1 Proof of Theorem 8
@@
LEMMA B.1.
Consider the matrix $H$ in (3.1). There is an absolute constant $C>0$ such that for any $D>0$, if $p\geq\frac{K^{2}e^{2(D+2)}(\log n)^{1+\frac{2}{\alpha}}}{n}$, with probability at least $1-2n^{-D}$,
$\|\widetilde{H}\|\leq C(1+\sqrt{D})\sqrt{np}.$ (B.1)
@@
Since the entries of $\widetilde{H}$ satisfy $\mathbb{E}|\widetilde{H}_{ij}|^{r}\leq K^{r}r^{\frac{r}{\alpha}}$ for all $r\geq 1$, applying Markov’s inequality to $|\widetilde{H}_{ij}|^{r}$ with $r=\log n$ and a union bound yields
$\mathbb{P}\left(|\widetilde{H}_{ij}|>e^{D+2}K(\log n)^{\frac{1}{\alpha}}\right)\leq n^{-2-D},$ (B.2)
$\mathbb{P}\left(\|\widetilde{H}\|_{\infty}>e^{D+2}K(\log n)^{\frac{1}{\alpha}}\right)\leq n^{-D}.$ (B.3)
Denote $B:=e^{D+2}K(\log n)^{\frac{1}{\alpha}}$. Define $\check{H}_{ij}=\widetilde{H}_{ij}\mathbf{1}\{|\widetilde{H}_{ij}|\leq B\}$. Then $|\check{H}_{ij}-\mathbb{E}\check{H}_{ij}|\leq 2B$. Since $\xi_{ij}$ is centered and has variance 1, for all $i,j$,
$\mathbb{E}(\check{H}_{ij}-\mathbb{E}\check{H}_{ij})^{2}=\mathrm{Var}(\check{H}_{ij})\leq\mathrm{Var}(\widetilde{H}_{ij})\leq p.$
Since $\check{H}$ has independent bounded entries, [44, Theorem 4.8] implies
$\|\check{H}-\mathbb{E}\check{H}\|\leq C\left(\sqrt{np}+B\sqrt{\log n}+t\right)$
@@
$\|\mathbb{E}\check{H}\|$ $=n|\mathbb{E}\check{H}_{ij}|=n|\mathbb{E}\check{H}_{ij}-\mathbb{E}\widetilde{H}_{ij}|=n|\mathbb{E}\widetilde{H}_{ij}\mathbf{1}\{|\widetilde{H}_{ij}|>B\}|$ (B.4)
$\leq n\sqrt{\mathbb{E}\widetilde{H}_{ij}^{2}}\sqrt{\mathbb{P}(|\widetilde{H}_{ij}|>B)}\leq Cn\sqrt{p}n^{-1-D/2}\leq C\sqrt{np}.$ (B.5)
@@
$\|\widetilde{H}\|=\|\check{H}\|$ $\leq\|\check{H}-\mathbb{E}\check{H}\|+\|\mathbb{E}\check{H}\|$ (B.6)
$\leq C(1+\sqrt{D})(\sqrt{np}+Ke^{D+2}(\log n)^{\frac{1}{2}+\frac{1}{\alpha}})\leq 2C(1+\sqrt{D})\sqrt{np},$ (B.7)
where in the last inequality, we use the assumption that $p\geq\frac{K^{2}e^{2(D+2)}(\log n)^{1+2/\alpha}}{n}$. ∎
We work on the conclusion of Lemma B.1. This result implies that for each eigenvalue $\lambda_{i}$ of $H$, $z_{i}:=\lambda_{i}+\mathrm{i}\eta$, where $\eta=LC_{\alpha}^{2}K^{4}\frac{\log^{2}n}{n}$, lies in a spectral domain
@@
$\mathbb{P}\left(\max_{i,j}\left|G_{ij}(z)-m(z)\delta_{ij}\right|\leq\delta\right)\geq 1-n^{-D}.$ (B.8)
@@
Therefore, taking $\delta=1/2$ in (B.8) and using the bound $|m|\leq 1$, we have
$\frac{3}{2}\geq\operatorname{Im}G_{kk}(z_{i})=\sum_{j=1}^{n}\frac{\eta}{(\lambda_{j}-\lambda_{i})^{2}+\eta^{2}}|u_{j}(k)|^{2}\geq\frac{1}{\eta}|u_{i}(k)|^{2},$ (B.9)
@@
B.2 Proof of Theorem 9
For $\alpha\geq 2$, we provide a stronger spectral norm concentration bound. The proof is based on a general result from [18] for the spectral norm of inhomogeneous random matrices where each entry has a symmetric distribution. We apply a symmetrization argument to obtain the following bound:
LEMMA B.2.
Let $X$ be an $n\times n$ real symmetric matrix with $X_{ij}=b_{ij}\xi_{ij}$ and $X_{ii}=\sqrt{2}b_{ij}\xi_{ii}$. Here $b_{ij}\geq 0$ and $\xi_{ij}$ are independent real random variables for $i\geq j$ with $\mathbb{E}\xi_{ij}^{2p}\leq\mathbb{E}g^{2p}$ for all $p\in\mathbb{N}$, where $g\sim N(0,1)$. Let $\sigma^{2}=\max_{i}\sum_{j}b_{ij}^{2},\sigma_{*}^{2}=\max_{ij}b_{ij}^{2}$. Then
$\mathbb{P}\left(\|X\|\geq 4\sqrt{\sigma^{2}+\sigma_{*}^{2}}+\sigma_{*}t\right)\leq 2ne^{-\frac{t^{2}}{16}}.$ (B.10)
Proof.
Let $X^{\prime}$ be a symmetric random matrix such that $X^{\prime}_{ij}=\varepsilon_{ij}X_{ij}$, where $\varepsilon_{ij}$ are i.i.d. Rademacher random variables. Then $X^{\prime}_{ij}$ has a symmetric distribution. From [59, Exercise 6.4.5] for any $t\geq 0$, $\mathbb{E}e^{t\|X\|}\leq\mathbb{E}e^{2t\|X^{\prime}\|}$. From [18, Lemma 3.17], assuming $\sigma_{*}=1$, and let $d=\lceil\sigma^{2}\rceil,$ we have $\mathbb{E}\operatorname{tr}e^{tX^{\prime}}\leq e^{2\sqrt{d}t+t^{2}}$. Since $e^{t\|X^{\prime}\|}\leq n\operatorname{tr}e^{tX^{\prime}}+n\operatorname{tr}e^{-tX^{\prime}}$, we obtain $\mathbb{E}[e^{t\|X^{\prime}\|}]\leq 2ne^{2\sqrt{d}t+t^{2}}$. Then $\mathbb{E}e^{t\|X\|}\leq 2ne^{4\sqrt{d}t+4t^{2}}$. By Markov’s inequality, for any $\varepsilon\geq 0$,
$\mathbb{P}\left(\|X\|\geq 4\sqrt{\sigma^{2}+1}+\varepsilon\right)\leq 2ne^{-\frac{\varepsilon^{2}}{16}}.$ (B.11)
Applying the inequality above to $\frac{1}{\sigma_{*}}X^{\prime}$, we obtain (B.10). ∎
With Lemma B.2, we obtain the following spectral norm concentration result for $\widetilde{H}$.
LEMMA B.3.
Consider the matrix $H$ in (3.1). Assume $\alpha\geq 2$. For any $c>0$, assume $p\geq\frac{c\log n}{n}$. For any $D>0$, there is a constant $C>0$ depending on $c,D$ such that with probability at least $1-2n^{-D}$, $\|\widetilde{H}\|\leq CK\sqrt{np}$.
Proof.
Since $\alpha\geq 2$, from the moment assumption on $\xi_{ij}$ we have
$\mathbb{E}(\xi_{ij})^{2p}\leq K^{2p}(2p)^{\frac{2p}{\alpha}}\leq K^{2p}(2p)^{p}.$
Since $\mathbb{E}g^{2p}=(2p-1)!!=\frac{2^{p}}{\sqrt{\pi}}\Gamma(p+1/2)$, and $\Gamma(x)\sim\sqrt{2\pi x}(x/e)^{x}$ as $x\to\infty$, there is an absolute constant $C>1$ such that $\mathbb{E}\left(\frac{1}{CK}\xi_{ij}\right)^{2p}\leq\mathbb{E}g^{2p}$. Let $X$ be an $n\times n$ symmetric random matrix such that $X_{ij}=x_{ij}$ where $x_{ij}\sim\mathrm{Ber}(p)$. Conditioned on $X$, $\widetilde{H}$ is a symmetric random matrix with a variance profile given by $X$. Since $p\geq\frac{c\log n}{n}$, by Bernstein’s inequality, there is a constant $C_{2}>0$ depending on $D$ such that with probability $1-n^{-D}$,
$\max_{i}\sum_{j}X_{ij}\leq C_{2}np.$
Conditioned on the event :=$\mathcal{E}:=\{\max_{i}\sum_{j}X_{ij}\leq C_{2}np\}$, by choosing $C_{2}$ sufficiently large, we can apply Lemma B.2 to $\frac{1}{CK}\tilde{H}$ to get
$\mathbb{P}\left(\frac{1}{CK}\|\tilde{H}\|\geq 3C_{2}\sqrt{np}\mid\mathcal{E}\right)\leq n^{-D}.$ (B.12)
Therefore with probability at least $1-2n^{-D}$, $\|\widetilde{H}\|\leq C^{\prime}K\sqrt{np}$ for some constant $C^{\prime}$ depending only on $c,D$. ∎
The proof of Theorem 9 follows a similar approach to that of Theorem 8. The key difference lies in the application of improved moment estimates in Lemma 6.2 and Lemma 6.4. For brevity, we present only the essential estimates, omitting the full derivation.
First, we have the analog of Theorem 6 as the following local law that holds up to the supercritical regime for $p$. Recall that we denote $\beta=1-\frac{1}{\alpha}$.
THEOREM 12 (Local law for sparse Wigner matrix with sub-gaussian entries).
Consider the matrix $H$ in (3.1) and its Green function $G(z)$. Assume $\alpha\geq 2$. For any $D>0$ and $\delta,\varepsilon\in(0,1)$, there exist $L\equiv L(\delta,D,\varepsilon)$ and $n_{0}\equiv n(\delta,\alpha,K,D,\varepsilon)$ such that the following holds for $n\geq n_{0}$. Let $p\geq LC_{\alpha}^{2}K^{4}\frac{\log n}{n}$. For $z=E+\mathrm{i}\eta\in\mathbf{S}$ where $\eta\geq n^{-1+\tau}$ with $\tau\geq\varepsilon(\log n)^{-1+\frac{1}{2\beta}}$, we have
$\mathbb{P}\left(\max_{i,j}\left|G_{ij}(z)-m(z)\delta_{ij}\right|\leq\delta\right)\geq 1-n^{-D}.$ (B.13)
Recall $\phi$ from (7.6). The proof of Theorem 12 relies on the next key estimates analogous to Lemma 7.1. These are derived using Lemma 5.3, Lemma 6.2, and Lemma 6.4, following a similar approach to the proof of Lemma 7.1.
LEMMA B.4.
Assume $r\geq 2$ and $\alpha\geq 2$. Denote $\beta=1-\frac{1}{\alpha}$. We have
$\max_{i\neq j}\|\phi G_{ij}\|_{L_{r}}\leq 4C_{\alpha}K\max\left\{\frac{\sqrt{r}}{\sqrt{n\eta}},\frac{r}{\sqrt{np}}\min\left\{1,\frac{1}{[\log(\frac{2r\eta}{p})_{+}]^{\beta}}\right\}\right\}:=\zeta_{1},$
$\max_{i,j\neq k}\|\phi(G_{ij}-G_{ij}^{(k)})\|_{L_{r}}\leq 4C_{\alpha}K\max\left\{\frac{\sqrt{r}}{\sqrt{n\eta}},\frac{r}{\sqrt{np}}\min\left\{1,\frac{1}{[\log(\frac{2r\eta}{p})_{+}]^{\beta}}\right\}\right\}=\zeta_{1},$
$\max_{i}\|\phi Y_{i}G_{ii}\|_{L_{r}}\leq 6C_{\alpha}K^{2}\max\left\{\frac{r}{\sqrt{n\eta}},\frac{r^{2}}{np}\min\left\{1,\frac{1}{[\log(\frac{r\sqrt{\eta}}{p\sqrt{n}})_{+}]^{2\beta}}\right\}\right\}$
$\qquad\qquad\qquad\qquad+6C_{\alpha}K^{2}\max\left\{\frac{\sqrt{r}}{\sqrt{np}},\frac{r}{np}\right\}:=\zeta_{2}.$
Specifically, $\zeta_{1}$ and the first term of $\zeta_{2}$ are obtained using the improved moment bounds from Lemma 6.2 and Lemma 6.4, while the second term of $\zeta_{2}$ follows from Lemma 5.3.
With Lemma B.4, we can prove the analogue of Theorem 11, following the same line of its proof. Define the fundamental error parameter $\widetilde{\zeta}:=\max\{\sqrt{\zeta_{1}},\sqrt{\zeta_{2}}\}$ where $\zeta_{1},\zeta_{2}$ are given in Lemma B.4.
PROPOSITION B.1.
Under the same assumption as in Theorem 12. Let $z\in\mathbf{S}$, $r\in\mathbb{N}$ and $p\in(0,1)$. Assume $n\geq 10$ and $r\geq 2$. For any $t\geq 1$ such that $20t\widetilde{\zeta}\leq 1$, we have
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z)-m(z)\delta_{ij}|>12t\tilde{\zeta}\right)\leq 3n^{6}t^{-2r}.$
Since Theorem 12 is valid for $p\gtrsim\frac{\log^{2}n}{n}$ as established in Theorem 6, it suffices to prove Theorem 12 for the regime that $\log n\lesssim np\leq\log^{3}n$ (say). Now we apply Proposition B.1 with $r=\log n$, $t=\exp(\frac{D+7}{2})$ and $\eta=n^{-1+\tau}$ with $\tau=\varepsilon(\log n)^{-1+\frac{1}{2\beta}}$. Denote $g(D,\delta):=\max\{\frac{12^{2}t^{2}}{\delta^{2}},{20^{2}t^{2}}\}$. Note that $\frac{2r\eta}{p}>1$ and $\frac{r\sqrt{\eta}}{p\sqrt{n}}>1$ for sufficiently large $n$. It is straightforward to verify that as long as
$p\geq\frac{40C_{\alpha}^{2}K^{4}}{\varepsilon^{2\beta}}g(D,\delta)^{2}\frac{\log n}{n},$
we have $12t\widetilde{\zeta}\leq\delta$ and $20t\widetilde{\zeta}\leq 1$ in Proposition B.1 for $n$ sufficiently large. Theorem 12 follows immediately.
Consequently, Theorem 9 can be proved following the same approach used for Theorem 8, incorporating Theorem 12 and Lemma B.3. The details are omitted.
Appendix C Proof of Proposition 7.1